OpenStudy (anonymous):

derivative of (2x-4)/(5x+5)

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( \frac{ u }{ v } \right)=\frac{ vu'-uv' }{v^2 }\]

OpenStudy (anonymous):

i did the same on wolfram alfa denominantor is same but upper part is +6

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

thanks surjit, i got it now